OpenStudy (anonymous):

For the following equation, state the value of the discriminant and the nature of the solutions. 6x^2-9x-8=0 What is the value of the discriminant? 1. The equation has one real solution. 2. The equation has two real solutions. 3. The equation has two imaginary solutions

OpenStudy (mertsj):

b^2-4ac=81-4(6)(-8)=81+192=273 Discriminate is positive so there are 2 real solutions

OpenStudy (2bornot2b):

The discriminants are \[\pm\sqrt{81-[4\times 6\times(-8)]}\]

OpenStudy (2bornot2b):

2. The equation has two real solutions.